write an equation in point-slope of the line through points (7,-4) and (8,1)
Answer by stanbon(75887)

write an equation in point-slope of the
line through points (7,-4) and (8,1)
-------------
slope = (1--4)/(8-7) = 5
---------
Form: y = mx + b
----
Substitute and solve for "b";
1 = 5*8 + b
b = -39
-------
Equation:
y = 5x - 39
=================
Cheers,
Stan H.
===========